# hep-ph/***FTUAM-21-$x$ IFT-UAM/CSIC-21-96

### v1

@@ 1 Introduction
let us note that indices are raised with the $\bar{g}_{\mu\nu}$ metric, and there is no closed formula in the general case. The same thing happens with the determinant of the metric.
@@
with $l$ a null vector, $l^{2}=0$. In this case the Ricci tensor can be written as
@@
This means that when the deformation $h_{\mu\nu}$ obeys the Fierz-Pauli equation [5] (please remember that $h^{\lambda}_{\lambda}=0$) the deformed metric is Ricci flat.
@@ 2 The non null Kerr-Schild gauge
On the other hand, the determinant $g$, using the matrix determinant lemma [7, 6], is related to the background determinant $\bar{g}$
@@
It is easy to check that there is also a relationship between the determinant of the deformation $h_{\mu\nu}$ and the determinant of the background metric $\bar{g}_{\mu\nu}$,
@@
but this second possibility corresponds to a rescaling of the metric
@@
Forgetting about this trivial rescaling, we will concentrate on the case $h=0$. Ttis means that the matrix $h$ has got rank $r\leq 3$. Therefore, the condition (2) has three different solutions
  • Rank one
@@
  • Rank two. Higher rank is more restrictive.
    $h_{\mu\nu}=\kappa A_{\mu}A_{\nu}+\kappa B_{\mu}B_{\nu}$ (29)
    with
    $A^{\lambda}A_{\lambda}=B^{\lambda}B_{\lambda}=-\frac{1}{2\kappa^{2}}\left(1+\frac{1}{\xi}\right)$
    $A^{\lambda}B_{\lambda}=0$ (30)
    this means that the vectors AA and $B$ are orthogonal and equally normalized. They cannot be both null. This condition could be softened by allowing two different $\xi_{A}\neq\xi_{B}$. There is a two-dimensional subspace of zero modes
    $A_{\mu}n_{i}^{\mu}=B_{\mu}n_{i}^{\mu}=0\quad i=1,2$ (31)
  • Rank three
    $h_{\mu\nu}=\kappa A_{\mu}A_{\nu}+\kappa B_{\mu}B_{\nu}+\kappa C_{\mu}C_{\nu}$ (32)
    where the vectors A, B and C are independent and non-null, in fact
    $A^{\lambda}A_{\lambda}=B^{\lambda}B_{\lambda}=C^{\lambda}C_{\lambda}=-\frac{1}{3\kappa^{2}}\left(1+\frac{1}{\xi}\right)$
    $A^{\lambda}B_{\lambda}=A^{\lambda}C_{\lambda}=B^{\lambda}C_{\lambda}=0$ (33)
    and there is now only a one-dimensional subspace of zero modes
    $A_{\mu}n^{\mu}=B_{\mu}n^{\mu}=C_{\mu}n^{\mu}=0$ (34)
From now on, we shall concentrate for simplicity in the rank one case. We will study the expression of the Ricci tensor, and its properties when applying the condition (25), concluding with a illustrative example. To begin with, the connection of the metric (16) is
@@
$\{R_{\mu\nu}=0\}\neq\{\bar{R}_{\mu\nu}=0\}$ (40)
We shall present now an explicit example of this fact
Start with a metric of Euclidean signature
@@
where the function $h_{2}[t]$ has the property $h_{2}^{\prime\prime}[t]=0$ and $\chi\equiv 1+\frac{1}{\xi}$. Then this deformation $h_{\mu\nu}$ obeys the Fierz-Pauli equation
@@ 3 The general deformation that obeys $R_{\mu\nu}=\bar{R}_{\mu\nu}$
Taking the trace, that is, multiplying by $\bar{g}^{\mu\nu}$
@@
This means that to have $R=\bar{R}$, we have three options, either $\xi=-1$, which corresponds with the null Kerr-Schild or else
$A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=0$ (52)
or the third possibility,
$A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=\phi A_{\beta}\hskip 28.45274pt\text{with}\hskip 28.45274pt\bar{\nabla}_{\alpha}A^{\alpha}=0$ (53)
@@
3.1 The first condition $A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=0$
@@
$\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\beta}A^{\alpha}-\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\alpha}A^{\beta}=0\Longrightarrow\bar{\nabla}_{\mu}A_{\nu}=\bar{\nabla}_{\nu}A_{\mu}$ (54)
This leads to
@@
3.2 The second alternative $A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=\phi A_{\beta}$ and $\bar{\nabla}_{\mu}A^{\mu}=0$
In this case,(49), implies
$\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\beta}A^{\alpha}-\bar{\nabla}_{\beta}A^{\alpha}\bar{\nabla}_{\alpha}A^{\beta}=\phi^{2}$ (56)
and a bit of algebra leads to
$R_{\mu\nu}=\bar{R}_{\mu\nu}-(1+\xi)\frac{\kappa^{2}}{2}\left[A^{\lambda}\bar{\nabla}_{\lambda}\left(\bar{\nabla}_{\mu}A_{\nu}+\bar{\nabla}_{\nu}A_{\mu}\right)]-\bar{\nabla}_{\lambda}A_{\mu}\bar{\nabla}^{\lambda}A_{\nu}+\bar{\nabla}_{\mu}A^{\lambda}\bar{\nabla}_{\nu}A_{\lambda}\right]-\xi\kappa^{4}\phi^{2}A_{\mu}A_{\nu}$
in conclusion, we get the desired result $R_{\mu\nu}=\bar{R}_{\mu\nu}$ only when $\phi=0$, in which case this case the second condition reduces to the previous one.

### v2

@@ 1 Introduction
let us note that indices are raised with the $\bar{g}_{\mu\nu}$ metric, there is no closed formula in the general case. The same thing happens with the determinant of the metric.
@@
with $l_{\alpha}$ a null vector, $l^{2}=0$. In this case the Ricci tensor can be written as
@@
This means that when the deformation $h_{\mu\nu}$ obeys the Fierz-Pauli equation [5] (please remember that $h^{\lambda}_{\lambda}=0$) then
@@ 2 The non null Kerr-Schild gauge
this idea was initially proposed in the paper [6].
On the other hand, the determinant $g$, with the matrix determinant lemma [7], is related to the background determinant $\bar{g}$
@@
It is easy to check that there is also a relationship between the determinant of the deformation $h_{\mu\nu}$ and the determinant of the background metric $\bar{g}_{\mu\nu}$
@@
but this corresponds with a rescaling of the metric
@@
Forgetting about this trivial rescaling, we will concentrate on the case $h=0$. This means that the matrix $h$ has got rank $r\leq 3$. Therefore, the condition (2) has three different solutions. Let us concentrate in the simplest, namely $r=1$, which corresponds to
@@
Next we will study the expression of the Ricci tensor, its properties when applying the condition (25), and conclude with a illustrative example.
First, the connection of the metric (16) is
@@
$R_{\mu\nu}\neq\bar{R}_{\mu\nu}$ (34)
We shall present an explicit example of this fact in a moment. Start with the Euclidean metric
@@
where the function $h_{2}[t]$ has the property $h_{2}^{\prime\prime}[t]=0$ and $\chi\equiv 1+\frac{1}{\xi}$. This deformation $h_{\mu\nu}$ obeys the Fierz-Pauli equation
@@ 3 The general deformation that obeys $R_{\mu\nu}=\bar{R}_{\mu\nu}$
taking the trace, that is, multiplying with $\bar{g}^{\mu\nu}$
@@
This means that in order to have $R=\bar{R}$, we have three options, either $\xi=-1$, which corresponds with the null Kerr-Schild or else
$\dot{A}_{\beta}\equiv A^{\alpha}\bar{\nabla}_{\alpha}A_{\beta}=0$ (46)
or finally the third possibility,
$\dot{A}_{\beta}=\phi A_{\beta}\hskip 28.45274pt\text{with}\hskip 28.45274pt\bar{\nabla}_{\alpha}A^{\alpha}=0$ (47)
@@
3.1 The first condition $\dot{A}_{\beta}=0$ (vanishing acceleration).
@@
$\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\beta}A^{\alpha}-\bar{\nabla}_{\beta}A_{\alpha}\bar{\nabla}^{\alpha}A^{\beta}=0$ (48)
This means that either
$\bar{\nabla}_{\beta}A_{\alpha}=0$ (49)
(in which case the deformed spacetime is automatically Ricci flat) or else
$\bar{\nabla}_{\mu}A_{\nu}=\bar{\nabla}_{\nu}A_{\mu}$ (50)
which leads to
@@
3.2 The second alternative $\dot{A}_{\beta}=\phi A_{\beta}$ plus $\bar{\nabla}_{\mu}A^{\mu}=0$
This is even simpler. Given the fact that $\overline{\nabla}_{\mu}A^{2}=0$ the acceleration is orthogonal to the deformation
$\dot{A}_{\mu}A^{\mu}=0$ (52)
Then $\phi=0$ for any non-null deformation, and we are back on the previous case.
@@ 4 Conclusions
Given a Ricci flat background seed metric any geodesic deformation will lead to a different spacetime, also Ricci flat. We are planning a systematic study of these in a future work.